OpenStudy (anonymous):

find the area of an equilateral triangle with an apothem 3.5 meters long and each side 12 meters long.

OpenStudy (angela210793):

A=b*h/2

OpenStudy (anonymous):

1}126 m 2}84 m 3}63 m 4}42 m

OpenStudy (anonymous):

\[\text{Area}=\frac{s^2n}{4 \text{Tan}\left[\frac{\pi }{n}\right]} \]Where n is the number of sides and s is the side length of a regular polygon. For this equilateral triangle, s=12 and n=3.\[\text{Area}=36 \sqrt{3}=62.4\]3} is the nearest value to the calculated area. http://www.mathopenref.com/polygonregulararea.html

OpenStudy (anonymous):

\[\text{Area}=a^2n \text{Tan}\left[\frac{\pi }{n}\right] \]Where n is the number of sides and a is the apothem of a regular polygon. The expression evaluated with a=3.5 and n=3 yields an area of 63.7 or 64 rounded.